There are three cousins who want to pool their money. Huey has d dollars. Louie has half of Huey's money. Dewey has 3 fewer dollars than Huey. If the sum of their money is $122, how much does each cousin have?

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52858)

.

After reading the problem, we write this

   H + L + D = 122  dollars.


Then we replace L with H/2 and replace D with (H-3) - - - the reasons why we do it are clear.


Then the equation takes form

    H + H/2 + (H-3) = 122.


Next you multiply everything by 2 and simplify step by step

    2H + H + 2H - 6 = 244

         5H         = 250

          H         = 250/5 = 50.


Thus we learned that Huey has 50 dollars.


From this point, complete the rest on your own.

Solved, with pretty full explanations.



Answer by greenestamps(13203)



The response from the other tutor shows one good standard formal algebraic solution.

If a formal algebraic solution is not required, the problem can be solved quickly and easily using logical reasoning and simple arithmetic.

Note that, even if you are asked for a formal algebraic solution, you can get good brain exercise and good practice in problem solving working the problem that way.

(1) Give Dewey $3 more, so that he and Huey now have the same amount, and the total is now $125.
(2) Louie now has half as much as each of the others, so the total is now 2.5 times the amount each of Huey and Dewey have.
(3) So the total $125 is 2.5 times the amount each of Huey and Dewey have; that means the amount each of them has is $125/2.5 = $50.
(4) Then take away the extra $3 you gave to Dewey.

ANSWERS: The amounts each of them has are
Huey: $50
Dewey: $50-$3 = $47
Louie: $50/2 = $25